OpenStudy (anonymous):

For all \[x >0\]and\[y >0\] the radical expression \[\frac{ \sqrt{x} }{ 5\sqrt{x}-\sqrt{y}}\] is equivalent to what?

OpenStudy (anonymous):

You multiply above and below by \[5\sqrt x+\sqrt y\]

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

So you remove the square root from the denominator

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so would it be \[\frac{ 5x+\sqrt{xy} }{ 25x-y }\]

OpenStudy (anonymous):

Exactly !

OpenStudy (anonymous):

alright thanks :) Would it be the same for all questions like this? Where I want to get rid of the radicals in the denominator?

OpenStudy (anonymous):

most of them ! ;)

OpenStudy (anonymous):

thnx a lot i appreciate it